whitemonsterbunny17 (whitemonsterbunny17):

Check my answer, please? ^.^ Thanks! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

Sketch the polygon described. If no such polygon exists, write "not possible". A regular polygon, one of whose angles has measure of 130. I believe it is \(\bf\color{red}{not~possible}\).

Nnesha (nnesha):

yes that is right

OpenStudy (jdoe0001):

hmmm what makes you think is not possible?

whitemonsterbunny17 (whitemonsterbunny17):

Because the interior measure cannot exceed 180, right?

OpenStudy (jdoe0001):

btw... .it doesn't say if it's an internal or external angle we're assuming is an internal angle I gather, but it really doesn't say

whitemonsterbunny17 (whitemonsterbunny17):

And I suppose so too, but the interior cannot have a measure greater that 108. right?

OpenStudy (jdoe0001):

hmmm you mean the SUM of interior angles in a TRIANGLE, yes... they sum up to 180 but that's just a triangle, which can be a regular polygon of 3 angles

whitemonsterbunny17 (whitemonsterbunny17):

Oh .-.

Nnesha (nnesha):

well i thin interior angle can have measure greater than 108 yes they can

OpenStudy (jdoe0001):

welll.... taking an assumption is an internal angle it's referring to sum of all internal angles in a polygon sum up to sides * angle = 180( sides - 2) so.... say we know the internal angles are 130 so angle = 130 sides = n n * 130 = 180( n -2) <---- if you solve for "n" to see how many sides it's.... what does it give you?

whitemonsterbunny17 (whitemonsterbunny17):

I have no idea

OpenStudy (jdoe0001):

if it were an external angle sum of all external angles sum up to 360 or sides * angle = 360 angle = 130 sides = n n * 130 = 360 <---- solving for "n"... how many sides does it have?

OpenStudy (jdoe0001):

well... is just a plain linear simplification.. surely you've covered that already, or you've been expected to

whitemonsterbunny17 (whitemonsterbunny17):

No, I have no idea what that is. .-. I know the formula to find how many sides, and I was alright at it, but now it confuses me.

OpenStudy (jdoe0001):

well... to find the angles... is the same as to get the sum of all angles really the equation can get either

OpenStudy (jdoe0001):

to find the sides rather...is the same equation as the sum of all angles.... the ones above ^

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so would I take 180-130 then take that number and divide 130 by it to find the number of sides?

OpenStudy (jdoe0001):

to find the sides assuming is 130 internal angle? sides * angle = 180( sides - 2) so.... say we know the internal angles are 130 so angle = 130 sides = n n * 130 = 180( n -2) <---- if you solve for "n" to see how many sides it's.... what does it give you?

whitemonsterbunny17 (whitemonsterbunny17):

Yeah, I know the formula, but I don't know how to get rid of the n's and still get a correct answer. Meaning I don't remember how to solve the algebra. .-.

OpenStudy (jdoe0001):

hmm you may want to brush up your linear simplification first then

whitemonsterbunny17 (whitemonsterbunny17):

Okay, but how? o.O

whitemonsterbunny17 (whitemonsterbunny17):

(And sorry, I know I'm a lot of work. I have a terrible memory when it comes to math stuff.)

whitemonsterbunny17 (whitemonsterbunny17):

Help? @sleepyjess .-.

OpenStudy (sleepyjess):

Bunny, here is a link i found http://www.algebra.com/algebra/homework/Triangles.faq.question.799486.html hopefully it helps